OpenStudy (anonymous):

what is the equation of the line that passes through (-9,12) and is perpendicular to the line whose equation is y=1/3*x+6? y=-1/3*x+15 y=-3*x-15 y=-1/3*x-13 y=-3*x+27

OpenStudy (texaschic101):

first, can you tell me what the slope of y = 1/3x + 6 is ??

OpenStudy (anonymous):

no idk

OpenStudy (khalilforthewin):

lord jesus

OpenStudy (khalilforthewin):

please bless this beautiful girl in hopes that she gets the right answer...amen

OpenStudy (texaschic101):

if not, I can show you how to find the slope....in y = mx + b, the slope can be found in the m position....so what is the slope of the original equation ?

OpenStudy (texaschic101):

y = mx + b y = (1/3)x + b do you see that the slope is 1/3 ?

OpenStudy (texaschic101):

we needed to find the slope because a negative reciprocal slope is needed for a perpendicular line. To find the negative reciprocal of 1/3, we just flip the slope and change the sign....so the new slope for out perpendicular line is -3/1....see how I flipped the slope and changed the sign ....any questions so far ?

OpenStudy (anonymous):

so its a or c

OpenStudy (texaschic101):

now we will use y = mx + b slope (m) = -3/1 (-9,12)...x = -9 and y = 12 now we sub and since we already know slope(m) and we have our points x and y, we just need to find b, the y intercept. so we sub y = mx + b 12 = -3/1(-9) + b 12 = 27 + b 12 - 27 = b - 15 = b so the perpendicular equation is : y = -3x - 15

OpenStudy (texaschic101):

any questions at all ??

OpenStudy (texaschic101):

well....if you have any, just ask